VIBRATIONAL ENERGY DISPOSAL AND RELATIVE REACTIVITY FOR THE VIBRATIONALLY DRIVEN REACTION OF CH_{3}D WITH Cl IN THE C-H C-D COMBINATION REGION

\underline{CHRISTOPHER ANNESLEY}, ROBERT J. HOLIDAY, CHAN HO KWON, F. FLEMING CRIM, Department of Chemistry, University of Wisconsin-Madison, Madison, Wisconsin 53706.

Product state distributions and action spectra for the reaction of vibrationally state selected CH_{3}D with Cl, CH_{3}D(ν) + Cl ⟶ (CH_{2}D + HCl) or (CH_{3} + DCl), reveal the extent to which vibrational energy initially deposited in the reactant survives in the CH_{2}D and CH_{3} products as well the reactivity of different excited vibrational states. There have been several studies of various deuterated methanes with different bonds excited. Most recently, our lab studied the C-H stretch overtone in the 6000 cm^{−1}region, where we found further support for the spectator model. We found that one C-H oscillator survived into the products from states that are composed of 2 separate C-H oscillators as opposed to an overtone in one oscillator. We are studying the lower energy 5000 cm^{−1}to 5500 cm^{−1}region, where we can excite combination bands that involve C-H and C-D motion to determine the relative effect on reactivity of each vibrational mode in this region.